How do I solve this? Please show detailed workings if possible, much thanks.
The function of f and g are given by f(x)=3^x-x and g(x)=9x+1
Find an equation of the line passing through the points (1,19) and (-1,-3)

Answer by Alan3354(69443)

The function of f and g are given by f(x)=3^x-x and g(x)=9x+1
Thx for letting us know.
====================
Find an equation of the line passing through the points (1,19) and (-1,-3)
Step 1, find the slope m
m = diffy/diffx
---
Step 2, use y-y1 = m*(x-x1) where (x1,y1) is either point.